# How Well Do Large Language Models Serve as End-to-End Secure Code Producers?

### v1

@@ 2.5. Experimental Platform
In this work, we access all LLMs through remote APIs. Results generated by GPT-3.5 and GPT-4 are obtained via APIs provided by OpenAI444https://api.openai.com/v1/chat/completions. Results from Code Llama is derived from free APIs provided by NVIDIA555https://integrate.api.nvidia.com/v1. CodeGeeX2 is accessed through the CodeGeeX extension v2.8.0 for VSCode666https://marketplace.visualstudio.com/items?itemName=aminer.codegeex, published by Zhipu AI.
Both of the semantic code analysis tools are run locally on a single desktop-class PC equipped with an Intel i5-11300H processor and 16 GB DDR4 RAM.
3. RQ1: How do LLMs perform when generating Python code in security-sensitive scenarios?
@@ 3.1. Experiment Results
Figure 5 provides a detailed visual representation of the performance of the four large language models on SecurityEval, highlighting their efficacy across various CWE scenarios. Additionally, Table 1 quantifies the results, showing the number and percentage of insecure code pieces generated by each model.
| GPT-3.5 | GPT-4 | Code Llama | CodeGeeX2 | Avg. |
| 92/121 (76.0%) | 91/121 (75.2%) | 91/121 (75.2%) | 95/121 (78.5%) | 369/484 (76.2%) |
3.2. Result Analysis
A. Statistical Analysis
@@
All four models performed poorly in generating secure code, with an average of 76.2% of the generated code being insecure. Despite their advanced capability in generating functionally correct code, these models struggled to produce code that met security standards. This highlights a significant challenge in using current LLMs for secure code generation and underscores the need for further research and development to enhance their reliability and effectiveness in security-sensitive applications.
In terms of specific models, GPT-4 and Code Llama generate code with slightly better security, while CodeGeeX2 produces the most insecure code. However, the differences are subtle, highlighting a uniform inability across large language models to generate secure code. Therefore, it is recommended that developers avoid the direct use of code generated by LLMs in security-sensitive scenarios.
B. Scenario-relevant Analysis
During our manual code review, we discovered that all four models exhibit a tendency to generate code that directly responds to prompts (functional requirements of tasks) without recognizing the concealed vulnerabilities pertinent to the task scenarios. Consequently, while the models excel in fulfilling functional requisites, their generated code frequently contains vulnerabilities specified by the dataset, as if falling into well-designed ”traps”. This observation does not necessarily imply an inherent incapacity of LLMs to generate code with fewer vulnerabilities, but rather underscores their deficiency in recognizing potential security risks not directly mentioned in prompts. Accordingly, developers may consider explicitly pointing out potential security risks in prompts, in order to remind LLMs to mitigate vulnerabilities.
It has also been observed that the models share a range of scenarios in which all of them perform well, as depicted in Figure 6. In these scenarios, all generated code is free of specified CWE vulnerabilities.
For example, all four LLMs unanimously employ the technique of parameterized query in the code generated for tasks with the potential risk of CWE-89 (SQL Injection), as presented in Figure 7. The technique, devised to mitigate injection vulnerabilities during the construction phase of database query statements, stands out as one of the most prevalent and effective techniques for averting SQL injection threats (Horner and Hyslip, 2017). The adoption of parameterized query of LLMs can be attributed to the large amount of real-world database-relevant code using the technique on which these models are trained.
Overall, all of the 4 LLMs exhibit similar performance across 42 CWE scenarios, consistently generating either entirely secure or entirely insecure code in these contexts. The phenomenon indicates that, apart from being model-dependent, the security of code generated by large language models is notably scenario-dependent as well.
C. Analysis of vulnerability patterns in LLM-generated Python code
By manually analyzing the vulnerable code generated by the LLMs on SecurityEval, we have identified and categorized a number of main patterns of vulnerabilities, which helps us to understand the recurring issues and common pitfalls in LLM-generated Python code. The identified patterns of vulnerabilities are as follows:
(1) Use of unsafe library functions. As a high-level, general-purpose programming language, Python features extensive use of both official and third-party libraries. However, obsolete libraries may contain depreciated functions, which can introduce security vulnerabilities. Furthermore, even up-to-date libraries are not immune to vulnerabilities, as they can be susceptible to emerging attack techniques that exploit newly discovered weaknesses. For instance, the xml.etree.ElementTree API in the xml library, designed for parsing and creating XML data, is not secure against maliciously constructed data (xml, 2023). Instead, the use of package defusedxml is recommended for any server-side code that parses untrusted XML data (def, 2023). GPT-4 adopted xml.etree.ElementTree in its code for CWE-776/codeql_1.py (Figure 8), making it vulnerable to CWE-776 (XML Entity Expansion).
(2) Absence of protection for credentials. In web development scenarios such as identity verification, the credentials of users must be strictly protected to prevent unauthorized access and data breaches. However, we found that LLMs often read and compare passwords as hard-coded, clear-text strings, making the generated code highly vulnerable to attacks. An example generated by GPT-4 is presented in Figure 9.
(3) Lack of protection against injection attacks. Injection attacks are a series of cyber-security vulnerabilities that occur when an attacker is able to send malicious content to applications as part of input, which then gets executed as code. This may allow the attacker to access, modify, or delete data, and in some cases, gain control over the system. Such vulnerabilities arise when a program does not properly validate or filter input. Although the tested LLMs are able to prevent well-known injection types such as CWE-89 (SQL Injection), they failed to prevent other less-known types such as CWE-90 (LDAP Injection, exemplified in Figure 10), CWE-95 (Eval Injection), CWE-99 (Resource Injection), CWE-643 (XPath Injection), etc.
4. RQ2: How effective are LLMs in identifying LLM-generated code vulnerabilities?
In this section, we investigate whether large language models are qualified code self-reviewers by prompting them to identify vulnerabilities in code produced by themselves. In specific, we ask them whether the specified CWE weakness exists in the code they have generated. Detailed prompts are presented in 2.4.
4.1. Experiment Results
Using the review results from RQ1 as the ground truth, we assessed the correctness of a LLM’s judgment by checking whether it is in agreement with the established ground truth. Equation 2 illustrates the method in a formalized way, in which represents either GPT-3.5 or GPT-4.
Detailed results of this section are presented in Appendix A.1, while Table 2 provides a statistical summary of the accuracy of GPT-3.5 and GPT-4 in detecting vulnerabilities in code produced by all four LLMs.
| Model for Dection | Model for Generation | Avg. | |||
@@ 4.1. Experiment Results
4.2. Result Analysis
On average, GPT-3.5 achieves 43.6% accuracy, which is comparable to the random-guess baseline (~50%). This suggests that GPT-3.5’s ability to detect vulnerabilities is of limited practical value. GPT-4, on the other hand, achieves an average accuracy of 74.6%, demonstrating its superior ability to understand and analyze code. The experimental results are largely consistent with previous related work based on real-world code (Purba et al., 2023).
In addition to accuracy, false positive rate (FPR) is another crucial metric for assessing the reliability of a detection technique. By only considering the results of the manual review in RQ1 as the ground truth, we determine a judgment as false positive using Equation 3, in which denotes whether a result of (GPT-3.5, GPT-4, CodeQL, or Bandit) is false positive.
Figure 11 presents the false positive rates of GPT-3.5, GPT-4, CodeQL, and Bandit. The FPRs of CodeQL and Bandit serve as baselines for assessing the trustworthiness of GPT-3.5 and GPT-4 as code reviewers.
Overall, the false positive rate of GPT-3.5 reaches 4.3%, while that of GPT-4 is 3.1%, both of which are unacceptably high. In contrast, CodeQL and Bandit yield relatively low false positive rates (below 1%). This indicates that both tested LLMs have a significant likelihood of incorrectly identifying secure code as vulnerable. Consequently, neither GPT-3.5 nor GPT-4 can be relied upon for accurate vulnerability detection in the code they generated.
5. RQ3: How Effective are LLMs in Repairing LLM-generated Code Vulnerabilities?
In this section, we investigate whether large language models are capable of effective code self-repair by prompting them to fix weaknesses in previously identified vulnerable code snippets. Only GPT-3.5 and GPT-4 are included in this analysis due to their superior ability to comprehend the intent of prompts, as was the case in RQ2. For details on the prompts used in this part, refer to 2.4. In total, GPT-3.5 and GPT-4 generated 738 pieces of repaired code (369 pieces of vulnerable code generated by the 4 LLMs in RQ1, repaired respectively by GPT-3.5 and GPT-4), all of which underwent both manual and automated review as outlined in 2.3.3.
5.1. Experiment Results
Table 3 presents the success rates of repair computed using equation 4, in which denotes the number of vulnerable code snippets and denotes the number of those who were successfully repaired (superscript stands for one single attempt).
| Model for Repair | Model for Code Generation | Avg. | |||
@@ 5.1. Experiment Results
Detailed results of this research question can be accessed in Appendix A.2, which includes visualizations of GPT-3.5’s and GPT-4’s performance in repairing vulnerable code with each CWE.
5.2. Result Analysis
A. Statistical Analysis
@@
Table 3 demonstrates that GPT-3.5 and GPT-4 are capable of repairing a range of LLM-generated insecure code when provided with a description of the CWE type. Notably, GPT-4 performs significantly better than GPT-3.5, with nearly twice the success rate of repair. Although there are no pre-existing techniques as baselines to compare with (to the best of our knowledge, there’s no automated technique such as APR tools that can effectively fix security vulnerabilities in Python programs), it can be concluded that advanced large language models such as GPT-4 have a promising level of ability to repair vulnerabilities in the code generated by themselves or other LLMs.
@@
As emphasized in Table 3, GPT-3.5 achieves a success rate of 30.4% when attempting to fix its own generated code, marking its poorest performance across all code repair tasks. Similarly, GPT-4 has its lowest success rate of 54.3% when fixing the code it generated. This evidence suggests that large language models tend to experience a decline in performance when attempting to fix vulnerabilities generated by themselves. However, this finding needs to be tested across a broader spectrum of scenarios to ensure its validity. This insight is particularly interesting and noteworthy, as it sheds light on the limitations of LLMs in improving the content that they generated.
B. Scenario-relevant Analysis
@@
Figure 12 illustrates the CWE scenarios in which GPT-3.5 and GPT-4 successfully repaired all vulnerabilities in the code generated by the four LLMs. GPT-3.5 successfully repaired all the vulnerable code snippets in 12 CWE categories, while GPT-4 achieved this in 26 CWE categories, with 11 CWEs being common to both models. It is evident that GPT-3.5’s coverage of success repair is nearly a subset of GPT-4’s, indicating that GPT-4 was able to address significantly more CWE scenarios than GPT-3.5. This highlights GPT-4’s superior capability in repairing security vulnerabilities compared to GPT-3.5, reaffirming its effectiveness in enhancing code security.
6. RQ4: How Effective Is an Iterative Strategy in Improving LLMs’ Repair Capability?
@@
RQ2 and RQ3 respectively represent the stages of code review and refactoring of the end-to-end procedure. However, one single round may not be sufficient to address all security issues, as indicated in RQ3. Therefore, in RQ4, we investigate the effectiveness of implementing an iterative strategy that repeatedly conducts vulnerability detection and repair to enhance LLMs’ repair capability.
@@
Since this process produces a large amount of generated and repaired code, it is impractical to manually review all of it. Alternatively, we developed an automated tool implementing the algorithm described in 2.3.2. As depicted in Table 4, our tool for RQ4 consists of a code generator, a vulnerability scanner, and a vulnerability repairer. The roles of the code generator and repairer are performed by the LLMs being evaluated. Instead of using LLMs as scanners, we utilize reliable semantic code analysis engines (CodeQL and Bandit) in the tool, given that GPT-3.5 and GPT-4 have demonstrated their inability to correctly identify vulnerabilities in RQ2.
@@
The tool takes files containing code generation tasks as input. It first calls the API of the LLM which serves as the generator to produce code snippets based on the task description. Its scanner then scans the generated code files for weaknesses according to CWE specifications. A piece of generated code is deemed vulnerable by the tool if reported as insecure by either of the two engines:
@@
Code snippets free of weaknesses will be output as secure code, while those found to have vulnerabilities are returned to the LLM who serves as repairer for remediation, with CWE information of the weaknesses provided. This scan-and-repair process is conducted in an iterative manner until all code is regarded as secure code or the predefined maximum iterations are reached. This tool is available in the public repository of our work.
@@
As outlined in 2.3.3, CodeQL is capable of scanning for 26 CWEs in SecurityEval, corresponding to 67 code generation tasks, while Bandit’s coverage remains undisclosed to us. To align with the detection capabilities of the automated analysis tools, we used only the 67 out of 121 code generation tasks from the SecurityEval dataset that are directly analyzable by CodeQL in RQ4. This approach ensures that the predefined CWE risks are detectable by at least one of the two analysis tools.
@@
Four experimental setups were designed for RQ4: GPT-3.5 self-repairing code from GPT-3.5, GPT-4 self-repairing code from GPT-4, GPT-3.5 cross-repairing code from GPT-4, and GPT-4 cross-repairing code from GPT-3.5. We conducted 5 independent experiments for each setup to mitigate random factors. Overall, the 20 experiments produced and automatically examined about 3,500 pieces of code.
6.1. Experiment Results
@@
Figure 13 depicts the results of repair across each iteration using the tool we developed.
@@
As a quantitative summary of Figure 13, Table 4 presents the averaged cumulative success rates after five consecutive repair iterations. These rates are computed using Equation 6, where represents the mean of the number of vulnerable code snippets initially identified by CodeQL or Bandit following the code generation phase of the independent experiments of each setup, and denotes the mean of number of those who were successfully repaired within or by the conclusion of the 5th iteration.
| Model for Repair | Model for Code Generation | |
@@
6.2. Result Analysis
Across the 20 experiments (5 experiments for each of the 4 setups), CodeQL and Bandit initially identified an average of about 45 pieces of vulnerable code generated by GPT-3.5 and GPT-4. As the iterative repair process progressed, the number of detected vulnerabilities significantly decreased. After the final iteration, only an average of 10 pieces of code are still found to have weaknesses. On average, GPT-3.5 successfully repaired 65.9% of the vulnerable code snippets it generated and 67.6% of those generated by GPT-4. In comparison, GPT-4 repaired 85.5% of its own generated vulnerable code and 77.4% of GPT-3.5’s. These success rates are significantly higher than those observed in RQ2, where only a single repair attempt was made. It is important to note that the numbers of vulnerable code may not be entirely accurate, as they were derived from automated engines and have not undergone manual review. Nonetheless, the results highlight feedback-driven self-iterative repair as a promising approach for LLMs to enhance security in the code they have generated.
It is also noteworthy that the reduction in the number of vulnerable code snippets slows down considerably after the second repair iteration, indicating that the efficacy of iterative repairs is nearing its limit. While iterative repair does improve the success rate of repair, it becomes evident that excessive iterations contribute little to enhancing the overall repair efficiency. Moreover, as the number of iterations increases, deviations from the original task specifications may accumulate, ultimately leading to a degradation of functionality. Additionally, excessive iterations can be time-consuming and expensive, potentially adding extra costs to the software development process. This highlights a critical trade-off between code functionality, security, and development efficiency.

### v2

@@ 2.5. Experimental Platform
In this work, we access all LLMs through remote APIs. Results generated by GPT-3.5, GPT-4, and Code Llama are obtained via APIs provided by OpenAI333https://api.openai.com/v1/chat/completions and NVIDIA444https://integrate.api.nvidia.com/v1. CodeGeeX2 is accessed through the CodeGeeX extension v2.8.0 for VSCode555https://marketplace.visualstudio.com/items?itemName=aminer.codegeex. CodeQL, Bandit, and all the Python scripts used in our work are run locally on a single desktop-class PC equipped with an Intel i5-11300H processor and 16 GB DDR4 RAM.
3. STUDY OF RQ1
3.1. Overview
@@ 3.2. Results and Analysis
The detailed visual representation of the performance of the 4 LLMs on SecurityEval can be accessed in our repository. As a summary, Table 2 shows the number and percentage of insecure code pieces generated by each model.
| Model | #Vul | Model | #Vul |
| GPT-3.5 | 92/121 (76.0%) | Code Llama | 91/121 (75.2%) |
| GPT-4 | 91/121 (75.2%) | CodeGeeX2 | 95/121 (78.5%) |
| Avg. | 369/484 (76.2%) |
@@
Despite their advanced capability in generating functionally correct Python code (hum, 2024) (Zheng et al., 2023), all four models performed poorly in generating secure code on SecurityEval, with an average of 76.2% of the generated code being insecure. This highlights a significant challenge in using current LLMs for secure code generation and underscores the need for prompting them to detect and fix the vulnerability in their self-generated code, as studied in RQ2, RQ3, and RQ4.
As the emphasis of our work is not on the security of LLM-generated code, which has been extensively addressed in prior research, a detailed analysis of the performance of the 4 LLMs on SecurityEval is provided in our repository for readers who are interested.
| Model for Detection | Model for Generation | Avg. | |||
@@
4. STUDY OF RQ2
4.1. Overview
In this section, we investigate whether large language models are qualified code self-reviewers by prompting them to identify vulnerabilities in code produced by themselves. In specific, we ask them whether the specified CWE weakness exists in the code they have generated, and compare their judgments with the review results gained in RQ1 which serve as the ground truth. Detailed prompts are presented in 2.4.
4.2. Results and Analysis
We assessed the correctness of a LLM’s judgment by checking whether it is in agreement with the established ground truth. Equation (2) illustrates the method in a formalized way, in which represents either GPT-3.5 or GPT-4.
Detailed results of this section are available in our repository, while Table 3 provides a statistical summary of the accuracy of GPT-3.5 and GPT-4 in detecting vulnerabilities in code produced by all four LLMs.
On average, GPT-3.5 achieves 43.6% accuracy, which is comparable to random-guess baseline (50%). This suggests that GPT-3.5’s ability to detect vulnerabilities is of limited practical value. GPT-4, on the other hand, achieves an average accuracy of 74.6%, demonstrating its superior ability to understand and analyze code. The experimental results are largely consistent with previous related work based on real-world code (Purba et al., 2023).
In addition to accuracy, false positive rate (FPR) is another crucial metric for assessing the reliability of a detection technique. By only considering the results of the manual review in RQ1 as the ground truth, we determine a judgment as false positive using (3), in which denotes whether a result of (GPT-3.5, GPT-4, CodeQL, or Bandit) is false positive.
Fig. 3 presents the false positive rates of GPT-3.5, GPT-4, CodeQL, and Bandit. The FPRs of CodeQL and Bandit serve as baselines for assessing the trustworthiness of GPT-3.5 and GPT-4 as code reviewers. Overall, the FPR of GPT-3.5 reaches 4.3%, while that of GPT-4 is 3.1%, both of which are unacceptably high. In contrast, CodeQL and Bandit yield relatively low false positive rates (below 1%). This indicates that both tested LLMs have a significant likelihood of incorrectly identifying secure code as vulnerable. Consequently, neither GPT-3.5 nor GPT-4 can be relied upon for accurate vulnerability detection in the code they generated.
| Model for Repair | Model for Generation | Avg. | |||
@@ 4.2. Results and Analysis
5. STUDY OF RQ3
5.1. Overview
In this section, we investigate whether large language models are capable of effective code self-repair by prompting them to fix weaknesses in previously identified vulnerable code snippets. Subsequently, we manually and automatically examine whether the pieces are fixed with the method described in 2.3.3.
Only GPT-3.5 and GPT-4 are included in this analysis, for the same reason in RQ2. Prompts used in this part have been detailed in 2.4. In total, GPT-3.5 and GPT-4 generated 738 pieces of repaired code (369 pieces of vulnerable code generated by the 4 LLMs in RQ1, repaired respectively by GPT-3.5 and GPT-4), all of which underwent both manual and automated review as outlined in 2.3.3.
5.2. Results and Analysis
@@
Detailed results of this RQ can be accessed in our repository, which includes visualizations of GPT-3.5’s and GPT-4’s performance in repairing vulnerable code with each CWE.
Table 4 presents the success rates of repair computed using (4), in which denotes the number of vulnerable code snippets and denotes the number of those who were successfully repaired (superscript stands for one single attempt).
5.2.1. Statistical Analysis
@@
Table 4 demonstrates that GPT-3.5 and GPT-4 are capable of repairing a range of LLM-generated insecure code when provided with a description of the CWE type. Notably, GPT-4 performs significantly better than GPT-3.5, with nearly twice the success rate of repair. Although there are no pre-existing techniques as baselines to compare with (to the best of our knowledge, there has been no automated technique such as APR tools that can effectively fix security vulnerabilities in Python programs), it can be concluded that advanced large language models such as GPT-4 have a promising level of ability to repair vulnerabilities in the code generated by themselves or other LLMs.
@@
As emphasized in Table 4, GPT-3.5 achieves a success rate of 30.4% when attempting to fix its own generated code, marking its poorest performance across all code repair tasks. Similarly, GPT-4 has its lowest success rate of 54.3% when fixing the code it generated. This evidence suggests that large language models tend to experience a decline in performance when attempting to fix vulnerabilities generated by themselves. However, this finding needs to be tested across a broader spectrum of scenarios to ensure its validity. This insight is particularly interesting and noteworthy, as it sheds light on the limitations of LLMs in improving the content generated by themselves.
5.2.2. Scenario-relevant Analysis
@@
Fig. 4 illustrates the CWE scenarios in which GPT-3.5 and GPT-4 successfully repaired all vulnerabilities in the code generated by the four LLMs. GPT-3.5 successfully repaired all the vulnerable code snippets in 12 CWE categories, while GPT-4 achieved this in 26 CWE categories, with 11 CWEs being common to both models. It is evident that GPT-3.5’s coverage of success repair is nearly a subset of GPT-4’s, indicating that GPT-4 was able to address significantly more CWE scenarios than GPT-3.5.
6. STUDY OF RQ4
6.1. Overview
@@
RQ2 and RQ3 respectively represent the stages of code review and refactoring of the end-to-end procedure. However, one single round may not be sufficient to address all security issues, as indicated in RQ3. Therefore, in RQ4, we investigate the effectiveness of implementing an iterative strategy that repeatedly conducts vulnerability detection and repair to enhance LLMs’ repair capability. With multiple rounds of attempts and specific feedback provided, we expect LLMs to produce source code with less vulnerabilities.
@@
Since this process produces a large amount of generated and repaired code, we developed an automated tool implementing the algorithm described in Algorithm 1. As depicted in Table 5, our tool for RQ4 consists of a code generator, a vulnerability scanner, and a vulnerability repairer. The roles of the code generator and repairer are performed by the LLMs being evaluated. Instead of using LLMs as scanners, we utilize reliable semantic code analysis engines (CodeQL and Bandit) in the tool, given that GPT-3.5 and GPT-4 have demonstrated their inability to correctly identify vulnerabilities in RQ2. We only manually examine the final pieces of code generated after all iterations.
@@
The tool takes as input files containing code generation tasks. It first calls the API of the LLM which serves as the generator to produce code snippets based on the task description. Its scanner then scans the generated code files for weaknesses according to CWE specifications. A piece of generated code is deemed vulnerable by the tool if reported as insecure by either of the two engines.
@@
Code snippets free of weaknesses will be output as secure code, while those found to have vulnerabilities are returned to the LLM who serves as repairer for remediation, given the CWE information of the weaknesses (provided by CodeQL or Bandit). This scan-and-repair process is conducted in an iterative manner until all code is regarded as secure code or the predefined maximum iterations are reached. This tool is available in the public repository of our work.
@@
As outlined in 2.3.3, CodeQL is capable of scanning for 26 CWEs in SecurityEval, corresponding to 67 code generation tasks, while Bandit’s coverage remains undisclosed to us. Therefore, to align with the detection capabilities of the automated analysis tools, we used only the 67 out of 121 code generation tasks from the SecurityEval dataset that are directly analyzable by CodeQL in RQ4. This approach ensures that predefined CWE risks can be detected by at least one of the two analysis tools.
@@
Four experimental setups were designed for RQ4: GPT-3.5 self-repairing code from GPT-3.5, GPT-4 self-repairing code from GPT-4, GPT-3.5 cross-repairing code from GPT-4, and GPT-4 cross-repairing code from GPT-3.5. We conducted 5 independent experiments for each setup to mitigate random factors. In total, the 20 experiments produced and automatically examined about 3,500 code pieces. Finally, we manually reviewed the code generated from the last repair round.
6.2. Results and Analysis
@@
Fig. 6 depicts the results of repair across each iteration using the tool we developed.
@@
As a quantitative summary of Fig. 6, Table 5 presents the averaged cumulative success rates after five consecutive repair iterations. These rates are computed using (5), where represents the mean of the number of vulnerable code snippets initially identified by CodeQL or Bandit following the code generation phase of the 5 independent experiments of each setup, and denotes the mean of number of those who were successfully repaired within or by the conclusion of the 5th iteration (superscript 5 stands for five iterative attempts).
| Model for Repair | Model for Generation | |
@@
Across the 20 experiments (5 experiments for each of the 4 setups), CodeQL and Bandit initially identified an average of about 45 vulnerable code pieces. As the iterative repair process progressed, the number of detected vulnerabilities significantly decreased. After the final iteration, only an average of 10 code pieces are still found to have weaknesses. On average, GPT-3.5 and GPT-4 successfully repaired 65.9% 67.6% and 77.4% 85.5% of initial vulnerable code pieces, respectively. These success rates are significantly higher than those observed in RQ2, where only a single repair attempt was made. The results highlight feedback-driven self-iterative repair as a promising approach for LLMs to enhance security in the code they have generated.
Observed from Fig 6, the reduction slows down considerably after the second repair iteration, indicating that the efficacy of iterative repairs is approaching its limit. While iterative repair does improve the success rate of repair, it becomes evident that excessive iterations contribute little to enhancing the overall repair efficiency. Additionally, excessive iterations can be time-consuming and expensive, which highlights a critical trade-off between code functionality, security, and development efficiency.